A piece of wire x feet in length is cut into exactly 6 pieces, each 2 feet 4 inches in length. What is the value of x ? (1 foot = 12 inches)
Answer by Theo(13342)

piece of wire is x feet in length.
cut into exactly 6 pieces.
each piece is 2 feet 4 inches in length.
what is the value of x?

equation is x = 6 * p where:

x is the length of wire.
p is the length of each piece of the wire.

length of each piece is 2 feet + 4 inches.

6 * (2 feet + 4 inches) = 6 * 2 feet + 6 * 4 inches = 12 feet + 24 inches.

24 inches = 2 feet.

length of wire = 12 feet + 2 feet = 14 feet.

14 feet divided by 6 = 2.333333333333 feet for each piece.

.333333333333 * 12 inches = 4 inches.

.33333333333 is the same as 1/3.

1/3 * 12 inches = 4 inches (same answer as above).

each piece is equal to 2 feet 4 inches.

answer is confirmed as correct.

you could also have converted everything to inches before starting.

6 * (2 feet 4 inches) is equal to 6 * (24 inches + 4 inches) is equal to 6 * 28 inches is equal to 168 inches.

168 inches / 12 = 14 feet.

length of wire equals 14 feet either way.